how many grams of potassium nitrate will you need to prepare 10L of a .75M solution
1
Expert's answer
2017-04-17T09:24:05-0400
СM = n/V n = m/M
M (KNO3) = 101 g/mol
n (KNO3) = СM x V = 0.75 x 10 = 7.5 mol
m (KNO3) = n x M = 7.5 x 101 = 757.5 g